# Analog model of Kerr black holes in scalar-tensor theories using binary Bose-Einstein condensates: Quasibound states and tachyonic instabilities

### v1

@@ Acknowledgements.
In addition to the analysis in frequency domain, we also study the time evolution of the small perturbations in the background of the presence of the spatial dependent mass term. By numerically integrating the equation (5) with the forth-order Runge-Kutta method, we focus on the response of the radial profile $H_{m}(r,t)$ at $r=2$ as a function of time. We consider the initial conditions of the Gaussian pulse
$H(r^{\ast},t=0)=\exp\left[-(r^{\ast}-\zeta^{\ast})^{2}/2\sigma^{2}\right],$ (S52)
$\partial_{t}H(r^{\ast},t=0)=0,$ (S53)
@@
Figure 6: The absolute value of time-domain profile with varying $\Omega_{0}$. The inset figure shows the fundamental QNM frequency with the corresponding value of $\Omega_{0}$. Other parameters are fixed as $m=1$, distance $r_{0}=15,$ and $\alpha=1/2$.
In Fig. 6, we compare three cases with $0.02\Omega_{0}=$ 0.0, $0.01$, and $0.1$. Even though the spectrum has significant change for small perturbations, such as $0.012\Omega_{0}=$ 0.01, we find that the response waveform in the time domain shows not much difference from the unperturbed one.

### v2

@@ Appendix C NUMERICALLY SOLVING THE TIME-DEPENDENT EQUATION IN (6) FOR $H_{m}(r,t)$
In addition to the analysis in frequency domain, we also study the time evolution of the small perturbations in the background of the presence of the spatial dependent mass term. By numerically integrating Eq. (6) with the forth-order Runge-Kutta method, we focus on the response of the radial profile $H_{m}(r,t)$ at $r=2$ as a function of time. We consider the initial conditions of the Gaussian pulse
$H(r^{\ast},t=0)=\exp\left[-(r^{\ast}-\zeta^{\ast})^{2}/2\sigma^{2}\right],$ (65)
$\partial_{t}H(r^{\ast},t=0)=0,$ (66)
@@
In Fig. 6, we compare three cases with $0.02\Omega_{0}=$ 0.0, $0.01$ and $0.1$. Even though the spectrum has significant change for small perturbations, such as $0.012\Omega_{0}=$ 0.01, we find that the response waveform in the time domain shows not much difference from the unperturbed one.